OpenStudy (anonymous):

Find an equation of the Tangent Line to the equation x^(2/3)+y^(2/3) = 4 at the point (-3sqrt(3), 1) the exponents in this really tripped me up!!! I ended up with the slope being 1/9^(1/18) I think this is wrong? can someone PLEASE help, I will award a medal! :) Please show how to solve this

OpenStudy (amistre64):

implicit .... works great

OpenStudy (amistre64):

x^(2/3)+y^(2/3) = 4 2/3 x^(-1/3) + 2/3 y^(-1/3) y' = 0 x^(-1/3) + y^(-1/3) y' = 0 y^(-1/3) y' = -x^(-1/3) y' = x^(-1/3) - --------- y^(-1/3) or simply: y' = - cbrt(y/x)

OpenStudy (amistre64):

-3sqrt(3), 1 y' = - cbrt(- 1/3sqrt(3))

OpenStudy (amistre64):

\[-\frac{(-1)^{1/3}}{(3(3)^{1/2})^{1/3}}\] \[-\frac{-1}{((27)^{1/2})^{1/3}}\] \[\frac{1}{(27)^{1/6}}=\frac{1}{3^{1/2}}\]

OpenStudy (amistre64):

27 = 3^3 ... so 3^3/6 = 3^1/2

OpenStudy (anonymous):

Thanks! I ultimately came up with (sqrt(3)/3)x +4